### v1

# ESTIMATING RESPIRATORY EFFORT FROM NOCTURNAL BREATHING SOUNDS FOR OBSTRUCTIVE SLEEP APNOEA SCREENING

Abstract
Obstructive sleep apnoea (OSA) is a prevalent condition with significant health consequences, yet many patients remain undiagnosed due to the complexity and cost of over-night polysomnography. Acoustic-based screening provides a scalable alternative, yet performance is limited by environmental noise and the lack of physiological context. Respiratory effort is a key signal used in clinical scoring of OSA events, but current approaches require additional contact sensors that reduce scalability and patient comfort. This paper presents the first study to estimate respiratory effort directly from nocturnal audio, enabling physiological context to be recovered from sound alone. We propose a latent-space fusion framework that integrates the estimated effort embeddings with acoustic features for OSA detection. Using a dataset of 157 nights from 103 participants recorded in home environments, our respiratory effort estimator achieves a concordance correlation coefficient of 0.48, capturing meaningful respiratory dynamics. Fusing effort and audio improves sensitivity and AUC over audio-only baselines, especially at low apnoea–hypopnoea index thresholds. The proposed approach requires only smartphone audio at test time, which enables sensor-free, scalable, and longitudinal OSA monitoring.
Index Terms—  Acoustic-based sleep apnoea detection, respiratory effort, obstructive sleep apnoea, nocturnal breathing sounds
1 Introduction
Obstructive sleep apnoea (OSA) is a prevalent form of sleep-disordered breathing (SDB) characterised by repeated episodes of partial or complete upper-airway obstruction during sleep. It affects approximately 16% of adults with higher prevalence in the older population [1, 2, 3]. Untreated OSA is associated with a range of health conditions, including cardiovascular disease, diabetes, and cognitive impairment [4]. Early detection and continuous tracking of OSA are therefore important, as timely intervention can reduce associated comorbidities and improve quality of life.
The diagnostic reference test for OSA is overnight polysomnography (PSG), which typically employs more than 12 sensors. Sleep clinicians manually score OSA events from these multimodal signals and compute the apnoea-hypopnoea index (AHI), defined as the number of events per hour of sleep. Although accurate, PSG is costly, labour-intensive, and restricted to specialised laboratories. The use of multiple sensors can disrupt sleep, and long waiting lists further delay diagnosis. Consequently, a large proportion of adults with OSA remain undiagnosed or untreated [5].
Recent evidence further highlights the limitations of one-night PSG assessments, with AHI demonstrating high night-to-night variability [6, 7]. This is clinically significant given its associations with other conditions. Longitudinal OSA monitoring provides a more reliable picture of habitual disease burden, yet current clinical pathways rarely enable frequent measurement.
Acoustic-based monitoring has been shown to be a promising non-invasive alternative. Snoring and breathing sounds carry physiological information about airway obstruction, and several studies have demonstrated feasibility of acoustic-based OSA screening. For example, Nakano et al. [8] reported strong performance for detecting mild/moderate/severe cases (AUC 0.99), using tracheal microphones in controlled hospital environments. In contrast, [9] investigated smartphone-based acoustic monitoring at home, which achieved competitive though lower AUCs of about 0.7 for mild and 0.8 for moderate/severe OSA. Similar declines in performance in home environments were reported in [10], who conducted a noise simulation and found acoustic-based OSA events detection is sensitive to specific noises including air conditioning and the human voice. To improve the robustness of audio-based methods, many studies introduce other physiological signals or demographic data as additional information [11, 12, 13]. Respiratory effort, in particular, has been shown to provide complementary information to acoustic features and significantly improve OSA screening [12]. However, current methods require additional sensors to capture effort, which limits scalability, increases cost, and reduces patient acceptability in home or longitudinal monitoring scenarios.
To overcome these barriers, we investigate whether respiratory effort cues can be inferred directly from audio. Our key hypothesis is that respiratory dynamics leave subtle imprints in snoring and breathing acoustics, and that explicitly modelling these latent cues can enhance the robustness of audio-based screening without introducing extra hardware requirements. The main contributions of this paper are: (i) We quantify the predictive relationship between snore acoustics and respiratory effort dynamics. (ii) We introduce a latent fusion strategy that integrates inferred respiratory effort with acoustic embeddings to regularise audio representations. (iii) We demonstrate that this approach improves OSA screening performance compared to audio-only baselines, while preserving the scalability and simplicity of purely acoustic methods.
The rest of the paper is organised as follows. Section 2 describes the dataset used in this study. Section 3 presents the proposed system for estimating respiratory effort from acoustics and integrating these embeddings for OSA detection. Section 4 describes the experimental setup, and results are reported and discussed in Section 5. Finally, Section 6 concludes the paper and outlines future directions.
Fig. 1: System diagram of the proposed latent fusion approach for acoustic-based sleep apnoea detection. A CNN-LSTM encoder extracts general acoustic embeddings and respiratory-effort-oriented latent representations from nocturnal audio. These are fused and fed into a classifier to predict apnoea/hypopnoea events.
2 Data
Experiments were conducted on a self-collected dataset [9] acquired using a Home Sleep Apnoea Test (HSAT) device (SOMNOtouch RESP). In parallel, nocturnal audio recordings were captured in real home environments using smartphones placed near the bed. The dataset consists of 157 nights from 103 participants, whose demographic and clinical characteristics are summarised in Table 1. The average nightly data duration is 7 hours and the total audio recording duration is just over 1,000 hours. Compared with other PSG datasets that include concurrent audio recordings, e.g. [14], this dataset provides a more balanced distribution across OSA severity levels.
The HSAT device records multiple physiological signals, including two respiratory effort channels derived from thoracic and abdominal movement sensors. In this study, we used the abdominal effort signal as the reference respiratory signal because it is less invasive and more comfortable for long-term monitoring. The abdominal effort signal was sampled at 32 Hz, while the audio recordings were sampled at 16 kHz. To synchronise the audio with the physiological signals, the snore channel from the HSAT device (sampled at 500 Hz), which correlates strongly with the audio signals, was used. Specifically, the audio was downsampled to 500 Hz and aligned with the snore channel by estimating the delay via cross-correlation [9]. After synchronisation, all model development and analysis were carried out using the original 16 kHz audio recordings.
Total Participants 103
Male 67 65%
Female 36 35%
Age (years) 45 ± 13 25 - 71
BMI (kg/m²) 31 ± 7 19 - 48
Total Nights 157
Night duration (hours) 7.0 ± 1.4 3.0 - 9.8
Healthy: AHI << 5 (nights) 16 10%
Mild: 5 $\leq$ AHI << 15 (nights) 59 38%
Moderate: 15 $\leq$ AHI << 30 (nights) 45 29%
Severe: AHI $\geq$ 30 (nights) 37 23%
Table 1: Demographic and clinical characteristics of the dataset used in this study.
3 Methods
All audio inputs were first segmented into overlapping 30-s windows with a 10-s shift. Each segment was transformed into 64-bin log-Mel filterbank features, with a 50-ms Hann window and a 20-ms frame shift. Each 30-s segment was therefore represented as a matrix of 1,500 frames $\times$ 64 frequency bins.
Fig. 1 shows the proposed two-step framework for OSA screening from audio recordings. In Step 1 (the blue part in Fig. 1), we trained an audio-to-effort estimator to infer respiratory effort directly from nocturnal audio. In Step 2 (the orange part in Fig. 1), the estimator was frozen and the final hidden state of its BiLSTM encoder was extracted as a respiratory embedding. This embedding was concatenated with the audio embedding from the audio-encoder, and the fused vector was passed through a fusion layer and a fully connected layer for classification. Importantly, no external sensors were required at test time.
3.1 Acoustic-based respiratory effort estimation
The proposed respiratory effort estimator consists of three main components: a convolutional neural network (CNN) feature extractor, an Long short-term memory (LSTM) encoder, and a decoder with interpolation. The CNN feature extractor uses smaller pooling kernels than the audio-only baseline to preserve temporal resolution, converting log-Mel spectrograms from (1500 $\times$ 64) to (187 $\times$ 128) (time frames $\times$ embedding dimensions). The LSTM encoder then processes these 187 frames to produce hidden vectors that capture temporal respiratory dynamics. For respiratory effort prediction, a linear decoder projects each hidden vector to a single value, generating a 187-point sequence.
The reference respiratory-effort signal is sampled at 32 Hz; each 30 s segment therefore has 960 points. As the predicted sequence contained only 187 points, it is subsequently interpolated to match the reference effort signal length. This approach reduces LSTM training complexity while maintaining prediction of amplitude and phase consistency through interpolation. Alternative approaches, such as pooling or linear projection, do not adequately reconstruct effort dynamics.
During training, losses are computed on these upsampled predictions. A key element of training is the use of the concordance correlation coefficient (CCC) [15] as the optimisation objective. CCC is defined as:
$\rho_{c}=\frac{2\sigma_{xy}}{\sigma_{x}^{2}+\sigma_{y}^{2}+(\mu_{x}-\mu_{y})^{2}},$ (1)
where $\mu_{x}$ and $\mu_{y}$ denote the means of the predicted and ground-truth signals, respectively, $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$ are their variances, and $\sigma_{xy}$ is their covariance. CCC ranges from 1-1 (no correlation) to $1$ (perfect correlation).
Unlike mean absolute error (MAE) and root mean square error (RMSE), CCC accounts for both correlation and bias, penalising mean and scale mismatches. We trained the model using $1-\rho_{c}$ as the loss function, and report CCC as the primary evaluation metric, with MAE and RMSE provided as complementary measures.
3.2 SDB events detection and AHI estimation
A separate audio-only CNN encoder is employed to extract audio embeddings for SDB event detection. As illustrated in Fig. 1, the proposed fusion framework integrates estimated respiratory embeddings with the audio embeddings to improve SDB event detection and AHI estimation.
Specifically, the respiratory embeddings are obtained by averaging the LSTM hidden states across time, producing a fixed-dimensional representation. This is concatenated with the audio embeddings from a baseline audio-only encoder, and the combined representation is used to predict the occurrence of apnoea and hypopnoea events. Finally, predicted events occurring in close succession are merged to form OSA episodes, from which night-level AHI is derived.
The AHI value is calculated as:
[…]
We employ a weighted binary cross-entropy loss to train the SDB event classifier, defined as:
$\mathcal{L}_{\text{BCE}}=-\frac{1}{N}\sum_{i=1}^{N}w_{i}\left[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\right],$ (3)
where $N$ is the total number of samples, $y_{i}$ is the ground-truth label, and $p_{i}$ is the predicted probability. The weight $w_{i}$ is defined as
$w_{i}=\frac{N}{2\times N_{c}},$ (4)
where $N_{c}$ is the number of samples in class $c$, used to address class imbalance between positive and negative events.
4 EXPERIMENTAL SETUP
4.1 Models for comparison
We first implemented an audio-only baseline using a similar CNN backbone and training procedure but without respiratory inputs, following [12]. The model comprised three convolutional blocks with 16, 32, and 64 filters, each followed by batch normalisation, ReLU activation, and max-pooling. The resulting feature maps were flattened, projected to a 512-dimensional vector, and then passed through a final fully connected layer with sigmoid activation to produce single logit.
We then evaluated our proposed latent-space fusion (LSF) model, which concatenated a frozen respiratory embedding (inferred from audio) with the audio-encoder output and passed the fused vector through a fusion layer and an MLP for classification.
For context, we also report an oracle results from prior work on the same dataset [12] that uses reference respiratory effort, representing an upper bound on performance.
4.2 Training setup
A 10-fold cross-validation scheme was employed for both respiratory effort estimation and AHI classification tasks. Data splits were performed at the subject level to ensure independence, with training, validation, and test sets following an 8:1:1 ratio. For each fold, we first trained the respiratory effort estimation module on the training set, then froze its parameters and used the extracted latent respiratory representations fused with audio CNN features for OSA classification training. Both effort estimation and classification performance were evaluated on the held-out test set of each fold.
4.3 Evaluation metrics
OSA event predictions were generated using a 30-s sliding window with a 10-s shift. Adjacent predicted OSA segments were merged to form contiguous OSA events, from which the night-level AHI was computed as described in Section 3.2.
The performance of respiratory effort signal estimator was evaluated at 30-s segment level with CCC, MAE and RMSE. Mean and standard deviation across all 10 folds are reported for each metric.
We used AHI cut-offs of 5, 15, and 30 events/h to stratify OSA severity groups (healthy, mild, moderate, severe). These thresholds are widely adopted for diagnosis and treatment of OSA [16]. An AHI $\geq$ 5 indicates OSA, and AHI $\geq$ 15 tends to lead to more arousals [17]. Untreated severe OSA (over 30) can lead to conditions such as hypertension and heart failure [18]. For each cut-offs, we evaluate the sensitivity, specificity and the area under the curve (AUC).
5 RESULTS AND Discussion
5.1 Respiratory effort prediction from audio
CCC $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
0.478 $\pm$ 0.133 1.053 $\pm$ 0.123 0.793 $\pm$ 0.092
Table 2: Evaluation metrics for respiratory signal prediction: Concordance Correlation Coefficient (CCC), Root Mean Square Error (RMSE), and Mean Absolute Error (MAE).
Table 2 reports the performance of respiratory effort prediction from nocturnal audio. Averaged across all folds, the proposed estimator achieved a CCC of 0.478, RMSE of 1.053, and MAE of 0.793. This suggest that audio carries information about underlying respiratory effort dynamics.
Fig. 2: Typical examples of predicted respiratory effort from audio (dashed line) compared to measured ground truth (solid line).
Fig. 2 illustrates a typical example of predicted respiratory effort from audio compared to the ground truth measured effort signal. In the first example, the estimated effort capture the overall trend of OSA events shown in the effort, with clear flat regions during the pause of snore sounds. Although the CCC correlation value is only around 0.48, the predicted effort successfully captures the physiological effort dynamics. The second example shows a close alignment with ground truth, while the third example highlights the situation of phase misalignment, where effort trends are reproduced but shifted in time. For segments lacking audible breathing, the predicted effort signals are nearly a flat line with little fluctuation. Overall, these results suggest that despite noise and variability, audio can provide a meaningful proxy for respiratory effort.
Related work using clean speech recordings on a smaller dataset (25 hours, 26 speakers) reported higher accuracy, with CCC of 0.6 and RMSE of 0.15 [19]. In contrast, our results reflect the challenges of a larger, noisier sleep sound dataset collected from real-world home environments using distant smartphone microphones. Higher MAE and RMSE may also stem from downsampling through pooling layers, which reduces temporal resolution of the effort signal. Additionally, temporal misalignment between audio and respiratory signals can lead to occasional low correlations, as illustrated in the bottom plot in Fig. 2.
5.2 OSA severity classification
AHI 5 15 30
157 nights 16 | 141 75 | 82 120 | 37
Audio only Sensitivity 0.86 0.81 0.78
Specificity 0.62 0.84 0.93
AUC 0.75 0.84 0.92
Latent Fusion – measured effort [12] Sensitivity 0.81 0.81 0.88
Specificity 0.82 0.74 0.87
AUC 0.87 0.86 0.93
Latent Fusion – estimated effort embeddings Sensitivity 0.88 0.83 0.78
Specificity 0.69 0.78 0.93
AUC 0.86 0.88 0.91
Table 3: OSA severity classification results at different AHI cut-off thresholds across models.
Table 3 lists the OSA severity classification results at different AHI cut-off thresholds across models. At the lowest AHI cut-off (5 events/h), the proposed LSF model with estimated respiratory effort embeddings achieves the highest sensitivity (0.88), slightly outperforming both the audio-only baseline (0.86) and the oracle system with measured effort (0.81). This suggests that incorporating estimated effort provides additional physiological context that is especially valuable for detecting mild OSA, where events are less pronounced and more difficult to identify from audio alone. However, this improvement in sensitivity comes at the expense of specificity, which drops compared to the oracle system (0.69 vs. 0.82), but still better than the audio-only baseline (0.62).
At the moderate AHI cut-offs (15 events/h), LSF with estimated effort again improves sensitivity over the baseline (0.83 vs. 0.81) while maintaining competitive specificity (0.78). The resulting AUC (0.88) is the highest across all systems, indicating that estimated effort embeddings contribute most strongly at clinically relevant mid-range severities where distinguishing between moderate and severe OSA is critical.
[…]
These results suggest that audio alone contains latent physiological cues that can support effective at-home OSA screening. The modest overall gains can be explained by several factors: (i) limited data and strong variability in smartphone recordings and real-world room acoustics, (ii) potential misalignment between audio and effort during zero-lag evaluation, and (iii) the coarseness of categorical AHI cut-offs compared to the continuous respiratory representation.
Overall, these results demonstrate that the proposed system, despite relying on noisy estimated effort signals inferred from audio, achieves AUC values close to those of models using measured effort. The key advantage of the proposed method is its scalability and non-invasiveness: at test time, only smartphone audio is required, which avoids additional sensors that may reduce patient acceptability or limit longitudinal monitoring.
6 Conclusion
In this study, we proposed a latent-space fusion framework that exploits respiratory effort embeddings inferred directly from nocturnal audio to improve OSA screening. Our experiments demonstrate that audio-inferred effort captures meaningful respiratory dynamics and, when fused with acoustic features, provides modest but consistent gains in sensitivity and AUC at clinically relevant AHI thresholds. Compared with prior work using measured respiratory effort [12], our approach achieves smaller improvements, reflecting the challenges of noise, temporal misalignment, and reduced precision in predicted effort signals.
Future work will explore simpler physiological targets, such as breathing-rate variability, to guide the audio-only classifier [20]. We also plan to improve effort prediction through domain-adapted snore/breathing foundation models, robust noise- and device-augmentation, and lag-tolerant training objectives. To employ respiratory acoustic foundation models [21, 22], additional fine-tuning to reduce the domain shift would be needed.